OpenStudy (anonymous):

Which is the interquartile range of the data set: 6, 8, 10, 12, 14?

OpenStudy (anonymous):

help me please

OpenStudy (amistre64):

a quartile is equal to 1/4 how many elements are in the set?

OpenStudy (anonymous):

5

OpenStudy (amistre64):

then the position 5/4 represents the first quartile and the position 3*5/4 represents the third quartile

OpenStudy (amistre64):

the interquartile range are the elements between those positions

OpenStudy (anonymous):

how did you get 3*5/4?

OpenStudy (amistre64):

6, 8, 10, 12, 14 |--------| IQR

OpenStudy (anonymous):

is there another way you can do this? i looked up on google and they said to find the median and all that junk

OpenStudy (amistre64):

a quartile represents 1/4 of a set. since there are 5 elements of the set, 5/4 is one quartile: the interquartile range is defined between q1 and q3 1*5/4 and 3*5/4

OpenStudy (amistre64):

the median is the position of q2; 2*5/4 = 2.xxx, so eounding up gives us the 3rd elements .... but thats not really needed for this

OpenStudy (amistre64):

*rounding up

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

1*5/4 = 1.25, since there is not 1.25 positional value; we rnd up to 2; the value in the 2nd position defines q1 3*5/4 = 3.75, since there is not 3.75 positional value; we rnd up to 4; the value in the 4th position defines q3 The IQR relates to the range from q1 to q3

OpenStudy (anonymous):

sigh idg it.. can yoou work it out?

OpenStudy (amistre64):

I just did, a few times even :)

OpenStudy (amistre64):

what part are you not understanding?

OpenStudy (anonymous):

oh right sorry.. oh wait i get it!! i just had to look at what you drew for me. i didnt see that lol.

OpenStudy (anonymous):

6, 8, 10, 12, 14 |--------| IQR

OpenStudy (anonymous):

you just put the two numbers together with the median because the other numbers had to be equal

OpenStudy (anonymous):

and thats how you got 3/5

OpenStudy (anonymous):

right? :) cus thats where i was confused @

OpenStudy (amistre64):

lets ignore the actual values of the elements: and just number them as: 1, 2, 3, 4, 5 this relates to a position within the set right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

the term quartile defines one quarter (1/4) of the set. does that make sense?

OpenStudy (anonymous):

yess

OpenStudy (amistre64):

then with 5 positions; divided into 4 parts defines the position of 1 quartile 5/4 = 1.25. is this position in our set?

OpenStudy (anonymous):

yes i think

OpenStudy (amistre64):

1, 2, 3, 4, 5 ; does not contain a positional value of 1.25.

OpenStudy (anonymous):

oh well i just thought you divided 5/4 lol

OpenStudy (anonymous):

well what do you do next then?

OpenStudy (amistre64):

i did, 5/4 = 1.25 which is what we need to use to define our positions by since 1.25 is not in our set, we want to go to the next whole value - which is 2

OpenStudy (anonymous):

ohh ok

OpenStudy (amistre64):

the 3rd quartile is determined in the same way: 3*5/4 = 3.75 is 3.75 in our set?

OpenStudy (anonymous):

i think

OpenStudy (amistre64):

1, 2, 3, 4, 5 does not contain the value 3.75, so we go to the next interger which is 4

OpenStudy (anonymous):

oh lol ok

OpenStudy (amistre64):

since the interquartile range (IQR) is the range from q1 to q3 ... 1, 2, 3, 4, 5 | IQR |

OpenStudy (amistre64):

relating this to the actual values in your set: 1 2 3 4 5 6, 8, 10, 12, 14 | IQR | 8 to 12

OpenStudy (amistre64):

this gets a slight twist IF our 1/4 is an integer, then we have to average some things ... but that wasnt the case for this :)

OpenStudy (anonymous):

okay i have to go soon. what do you do next?

OpenStudy (amistre64):

you write down the answer as 8 to 12 and move to the next question .....

OpenStudy (amistre64):

if there are options to pick, then you pick the one that says 8 to 12 ....

OpenStudy (anonymous):

ok my answer options are: a. 7 c. 6 b. 13 d. 10

OpenStudy (anonymous):

so it would be 6 or 10

OpenStudy (amistre64):

those are odd options; but i would say it would be 10, since that is the only one that is actually a part of the IQR: 8,10,12